OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 8. Find y when x = 2.

OpenStudy (anonymous):

Given: y = k/(2x) Plug in the given values: 8 = k/(2(3)) = k/6 Solve for k: k = (8)(6) = 48 Plug that back into the original equation: y = 48/(2x) = 24/x Let x = 2:

OpenStudy (anonymous):

y = 24/2 = 12

OpenStudy (anonymous):

sorry it cut me off there haha